OpenStudy (anonymous):

19k+18-22k=16-(3k-2) linear equation

OpenStudy (anonymous):

@izuru

OpenStudy (anonymous):

@Leong

OpenStudy (anonymous):

Are you trying to solve for "k"?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@raesal

OpenStudy (anonymous):

Get all the K terms on one side and all the non K terms on the other

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

19k+18-22k=16-(3k-2) -3k+18=16-3k+2 0=0k??? I think i wrong somewhere

OpenStudy (anonymous):